2. You have obtained the following information on a country’s population: 

Number of employed people who are looking for work every day: 420 

Total number of employed people: 1000 

Total number of unemployed people: 30 

Number of retired people: 200 

People over the age of 16: 2000 

a)Calculate the unemployment rate and labor-force-participation rate of this country

b)What do you understand by the Menu Cost of inflation? Explain with example.

3. In the year 2010, Burgerina's per capita GDP was 142.69 USD. Due to extraordinary economic policy  measures implemented by the government, the country witnessed a staggering growth of 10% for the last  11 years. What is Burgerina’s GDP in 2021?


1
Expert's answer
2021-11-09T10:45:38-0500

2.

(a)

Unemployment rate is obtained by dividing bthe number of unemployed persons by the number of persons in the labor force ( whether employed or unemployed) and multiplying that figure by 100.

Unemployment rate"=\frac{30}{100+30+200}\times100=2.439" %.

Participation rate is the percentage of the population that is either working or actively looking for work.

Labor force participation rate is obtained by dividing labor force by civilian non-institutional population and the multiplying by 100.

"=\frac{1000+30}{2000}\times100 =51.5"


(b)

Menu cost of inflation

Menu cost is the cost to a firm resulting from changing its prices. Menu costs are usually the result of inflation. Money loses value with inflation leading to a drop innthe purchasing power of an individual dollar.

For example, when raising prices, there are additional costs e.g. printing new menus, updating the website.

(3)

Find 10 percent of 142.69$.

"=\frac{142.69\times10}{100}=14.269"

Multiplying by 11 years:

"=14.269\times11=156.959"

GDP in 2021:

"142.69+156.959=299.649"